# SQL-Factory: A Multi-Agent Framework for High-Quality and Large-Scale SQL Generation

### v1

@@ Abstract.
Hight quality SQL corpus is essential for intelligent database. For example, Text-to-SQL requires SQL queries and correspond natural language questions as training samples. However, collecting such query corpus remains challenging in practice due to the high cost of manual annotation, which highlights the importance of automatic SQL generation. Despite recent advances, existing generation methods still face limitations in achieving both diversity and cost-effectiveness. Besides, many methods also treat all tables equally during generation, which overlooks schema complexity and leads to under-utilization of structurally rich tables. To address these issues, this paper proposes a multi-agent framework for high-quality and large-scale SQL generation, dubbed SQL-Factory. It decomposes the generation process into three collaborative teams: the Generation Team explores diverse query structures using large language models, the Expansion Team scales promising patterns via lightweight local models, and the Management Team adaptively schedules and evaluates generation based on schema coverage and real-time query quality. This modular framework ensures a balanced trade-off between diversity, scalability, and generation cost. We apply SQL-Factory to four widely used benchmarks and generate over 300,000 executable and broadly distributed SQL queries with less than \$200 API cost. Our generated queries achieve higher diversity compared to other methods, and extensive experiments demonstrate that the generated queries significantly improve the model performance in various downstream tasks.
@@ 1. Introduction
Insufficient query diversity. A lack of diversity means that the generated SQL queries tend to follow similar structures, use the same types of operators, and involve limited combinations of joined tables. Template-based approaches such as TPC-DS (Nambiar and Poess, 2006) and Join Order Benchmark (JOB) (Leis et al., 2018) typically rely on fixed syntactic forms, which restrict their ability to produce varied query types. Learning-based methods such as TreeGAN (Liu et al., 2018) and LearnedSQLGen (Zhang et al., 2022) exhibit strong tendencies to overfit to frequent query skeletons in training corpora, producing workloads clustered around recurrent table subsets and operator chains. As a result, these methods fail to capture the rich variety of queries that appear in real-world workloads. Improving query diversity would help cover more edge cases and make trained models more robust in practical database environments.
@@ 2.2. SQL Generation
SQL generation aims to automatically synthesize large volumes of semantically meaningful and structurally diverse SQL queries tailored to a specific database schema or system. This task plays an essential role in supporting downstream applications such as system benchmarking, robustness testing, and data-driven model training.
@@ 2.3. LLM-based Multi-Agent System
Large language models (LLMs) have demonstrated remarkable capabilities in natural language understanding and reasoning, enabling them to perform a wide range of complex tasks with human-like proficiency. In single-agent architectures, an LLM is typically equipped with modular components such as Profile, Memory, Planning, and Action (Xi et al., 2025), allowing it to decompose tasks, retrieve relevant information, and simulate multi-step workflows. This design has proven effective in various domains, including dialogue generation, retrieval-augmented generation, and structured data interaction (Wang et al., 2024; Modarressi et al., 2023; Hu et al., 2023).
@@
Building upon this foundation, recent research has extended LLM capabilities into collaborative multi-agent systems, where multiple specialized agents coordinate to complete tasks through dynamic communication and role-based reasoning. Unlike single-agent models that rely on a monolithic structure, multi-agent systems distribute cognitive responsibilities across distinct agents, each tailored for specific subgoals or expertise areas. This paradigm fosters more interpretable, scalable, and resilient behavior, particularly in settings that demand iterative planning, task delegation, and adaptive feedback handling. Multi-agent LLM systems have shown promise in domains such as software engineering (Qian et al., 2024; Chen et al., 2023), where agents representing roles like developer, tester, and reviewer to build and refine software artifacts. Similarly, in scientific experimentation (Zheng et al., 2023) and medical diagnostics (Tang et al., 2024), these systems enable complex reasoning chains by leveraging role specialization and interactive deliberation among agents.
@@
In the context of databases, LLM-based multi-agent approaches have recently emerged as a powerful solution for tasks requiring high modularity and semantic precision. For instance, in Text-to-SQL translation, multiple agents can collaboratively handle subtasks such as schema linking, query generation, and final SQL refinement, leading to more robust and accurate SQL outputs (Wang et al., 2025b; Xie et al., 2024). In the realm of database anomaly diagnosis, systems like D-Bot leverage large language models to automatically analyze and identify root causes of anomalies by extracting knowledge from diagnostic documents and employing structured reasoning techniques (Zhou et al., 2024). Additionally, ROMAS introduces a role-based multi-agent framework for database monitoring and planning, featuring agents assigned to specific roles (planner, monitor, and worker) collaboratively facilitate dynamic task allocation, self-monitoring, and adaptive re-planning in complex data environments (Huang et al., 2024). These examples collectively demonstrate the potential of multi-agent architectures to enhance reasoning modularity, improve decision accuracy, and integrate system-level feedback more effectively than traditional single-agent approaches. Despite these advances, the application of LLM-based multi-agent systems to large-scale SQL generation remains underexplored.
@@ 3. System Overview
In this section, we present an overview of SQL-Factory, a multi-agent framework designed to automatically synthesize large-scale, structurally diverse, and executable SQL query workloads in a cost-efficient and schema-aware method. Unlike prior methods that struggle to balance generation quality with scalability, SQL-Factory explicitly addresses three core challenges in SQL generation: ensuring sufficient structural diversity, reducing the cost of large-scale synthesis, and promoting balanced schema utilization by prioritizing structurally rich or underrepresented tables.
Figure 1 provides an overview of SQL-Factory, which is organized into three collaborative agent teams: (1) Management Team, (2) Generation Team, (3) Expansion Team, along with a centralized (4) SQL Pool. The Management Team comprises two agents: the Management Agent and the Critical Agent. The Generation Team includes the Table Selection Agent and the Generation Agent, while the Expansion Team is made up of the Seed Selection Agent and the Expansion Agent. These three teams collaborate synergistically to streamline the generation, expansion, and validation of SQL queries, with each team dedicated to a specialized role. The SQL Pool, implemented using a relational database such as MySQL, serves as the centralized repository for all queries.
Management Team. The Management Team governs the overall coordination of SQL synthesis by maintaining a dynamic balance between exploration (via the Generation Team) and exploitation (via the Expansion Team). It is composed of the Management Agent and the Critical Agent. The Critical Agent provides real-time evaluations of generated and expanded queries using two dedicated tools: an Execution Tool to verify semantic and syntactic correctness, and a Retrieval Tool to help assess diversity. Leveraging these evaluations, the Management Agent adaptively schedules the Generation and Expansion Teams. This dynamic adjustment allows SQL-Factory to self-correct in response to redundancy or quality degradation, ensuring an effective trade-off between structural exploration and pattern exploitation.
Generation Team. The Generation Team focuses on exploration by synthesizing diverse and structurally complex SQL queries from scratch using a powerful language model. It is guided by schema-aware planning, which prioritizes structurally rich and underutilized tables based on schema complexity and query distribution. This team generates high-diversity queries that incorporate advanced SQL constructs such as nested subqueries, joins, aggregations, and window functions. By generating queries from scratch, the Generation Team helps introduce novel structures and creates new patterns that guide subsequent expansion.
@@
Expansion Team. The Expansion Team focuses on exploitation by rapidly scaling the quantity of executable queries while preserving semantic richness. It reuses seed queries from the SQL Pool and employs a lightweight, cost-efficient model to synthesizes new variants through controlled augmentation. Unlike the Generation Team, which focuses on novelty, the Expansion Team concentrates on scaling promising patterns to achieve high query volume with minimal computational overhead, enabling the system to achieve both depth and breadth in SQL generation.
@@
In this section, we provide a detailed methodology for SQL-Factory by describing the internal design and interactions of its key components, including the Management Team, Generation Team and Expansion Teams.
@@ 4.1.1. Management Agent
(1) $S_{t+1}=\text{Agent}_{\text{M}}(S_{t},F_{t}),$
@@
Specifically, the system begins in the exploration phase, where the Generation Team synthesizes queries for underrepresented tables. After a fixed number of generation cycles, control shifts to the Expansion Team to exploit learned patterns. During this exploitation phase, the Management Agent continuously monitors evaluation information from Critical Agent. If query redundancy increases significantly or executability declines, it triggers a transition back to exploration. This alternating control mechanism allows SQL-Factory to incrementally grow a dataset that is both diverse and scalable, while avoiding mode collapse or overfitting to common patterns.
@@ 4.1.2. Critical Agent
Specifically, the Critical Agent comprises two modular components: the Execution Tool, which verifies query executability, and the Retrieval Tool, which using hybrid similarity to retrieve the most similar SQL queries. The evaluation process differs slightly depending on whether the query was produced by the Generation Team or the Expansion Team.
For SQL queries from the Generation Team, the powerful large model ensures a degree of novelty and exploratory coverage. Therefore, we focus primarily on syntactic and semantic validity. Given a set of queries $\mathcal{Q}=\{q_{t}^{1},q_{t}^{2},\dots,q_{t}^{n}\}$ at time step $t$, the Critical Agent call the Execution Tool $\text{Tool}_{\text{E}}$ to compile each query on the target database schema and reports a binary executability flag. The overall evaluation for this batch is defined as:
(2) $E_{t}=\text{Agent}_{\text{C}}(\mathcal{Q})=\frac{1}{n}\sum_{i=1}^{n}\text{Tool}_{\text{E}}(q_{t}^{i}),$
where $\text{Tool}_{\text{E}}(\cdot)\in\{0,1\}$ returns whether the query compiles successfully or not.
@@
By integrating these three level signals, this hybrid scoring function provides a more comprehensive assessment of query similarity. It is particularly effective in identifying subtle redundancies among expanded SQL queries, and helps ensure that only genuinely novel and diverse queries are admitted into the SQL Pool.
To systematically combine both executability and diversity assessments, the final evaluation for the Expansion Team is defined as:
(6) $E_{t}=\left[\frac{1}{n}\sum_{i=1}^{n}\text{Tool}_{\text{E}}(q_{t}^{i}),\;\frac{1}{n\cdot k}\sum_{i=1}^{n}\text{Sim}(q_{t}^{i},\;\text{Tool}_{\text{R}}(q_{t}^{i}))\right],$
where $\text{Tool}_{\text{R}}(q_{t}^{i})$ returns the query set of the top-$k$ nearest neighbors of $q_{t}^{i}$ using the hybrid similarity metric. The final evaluation vector $E_{t}$ is then passed to the Management Agent to support real-time scheduling decisions.
@@ 4.2.1. Table Selection Agent
(8) $\text{Agent}_{\text{T}}=\left\{\text{Schema}(T_{i})\mid T_{i}\in\mathcal{T}^{\prime},\ \mathcal{T}^{\prime}\subset\mathcal{T}\right\}$
@@ 4.2.2. Generation Agent
The agent is implemented using a powerful, instruction-tuned large language model, which is prompted with both schema information for $\mathcal{T}^{\prime}$ and high-level generation instructions, as shown in Figure 3. Rather than treating each table independently, the model considers the joint schema of all selected tables and is encouraged to synthesize complex queries that include multi-table joins, nested subqueries, aggregations, and window functions. Besides, to guide the model effectively, the schema information included in the prompt is carefully constructed to ensure both syntactic validity and semantic meaningfulness of the generated queries. For each selected table, we encode not only its column names and data types, but also relevant content-aware hints to reduce the likelihood of generating vacuous or logically invalid queries. Specifically, if a column is of an enumerated type (e.g., status, category), we provide up to five representative values from the domain to help the model infer plausible filter conditions or join constraints. If the column is of a numeric or date type (e.g., int, float, date), we include its minimum and maximum values observed in the database, offering statistical context for range-based predicates (e.g., price ¿ 100). For free-text or string columns, we sample up to five typical values as textual examples, which can guide the model to generate meaningful LIKE conditions or equality comparisons. These content-aware augmentations are inserted inline with the schema description in a structured format, forming an enriched prompt that provides the model with both structural knowledge and semantic grounding. This design significantly improves the quality and contextual relevance of the generated SQL queries, particularly for tables with sparse usage history or rich attribute semantics.
Finally, the input and output of the Generation Agent are summarized as:
@@ 4.3.1. Seed Selection Agent
To select appropriate seeds, the agent first determines a set of underutilized tables $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ by analyzing query distribution statistics via a shared statistical module $\text{Tool}_{\text{S}}$. It then search the SQL Pool to collect a small set of seed queries $\mathcal{Q}{\text{seed}}$, each of which must reference at least one table from $\mathcal{T}^{\prime}$. These queries are selected not only based on table inclusion but also for their structural validity and potential for lexical or syntactic variation. This approach allows the Expansion Agent to operate in a pattern-aware and data-driven manner, leveraging past structures to efficiently generate new and diverse SQL variants. Formally, the the final output of Seed Selection Agent is:
(10) $\mathcal{Q}_{\text{seed}}=\text{Agent}_{\text{S}}(\mathcal{T}^{\prime})=\text{Random}(\{q_{i}\mid\text{Tables}(q_{i})\cap\mathcal{T}^{\prime}\neq\emptyset\}).$
@@ 4.3.2. Expansion Agent
Figure 4. Expansion Prompt Template.
Specifically, this agent operates on the set of seed queries, which were selected based on their relevance to underrepresented tables. It uses a lightweight language model to derive multiple variants based on these seed queries, employing techniques such as structural rewriting, substructure recombination, and schema-guided composition. Besides, the agent is not limited to structural transformations but also introduces lexical and logical variations to ensure semantic diversity.
As shown in Figure 4, each prompt includes a small set of seed SQL queries and task-specific instructions. These instructions encourage the model to generate structurally diverse and executable variants, while inferring schema constraints implicitly from the provided queries. Formally, given a set of seed queries $\mathcal{Q}{\text{seed}}$, the Expansion Agent generates a corresponding set of SQL variants $\mathcal{Q}\text{exp}$ as follows:
@@
After expansion, the resulting queries $\mathcal{Q}_{\text{exp}}$ are passed to the Critical Agent. In this phase, queries are not only checked for executability but also evaluated for semantic redundancy using retrieval-based similarity assessment. Only queries that meet the criteria for both executability and novelty are retained and added to the SQL Pool.
Through this process, the Expansion Agent supports vertical dataset expansion. It strengthens the density of high-quality query patterns within the local neighborhood of schema space, ultimately improving the dataset’s expressiveness and utility in downstream tasks. While the Generation Agent explores new structural forms, the Expansion Agent intensifies local pattern regions, ensuring both diversity and volume.

### v2

@@ Abstract.
Hight quality SQL corpus is essential for intelligent database. For example, Text-to-SQL requires SQL queries and correspond natural language questions as training samples. However, collecting such query corpus remains challenging in practice due to the high cost of manual annotation, which highlights the importance of automatic SQL generation. Despite recent advances, existing generation methods still face limitations in achieving both diversity and cost-effectiveness. Besides, many methods also treat all tables equally, which overlooks schema complexity and leads to under-utilization of structurally rich tables. To address these issues, this paper proposes a multi-agent framework for high-quality and large-scale SQL generation, dubbed SQL-Factory. It decomposes the generation process into three collaborative teams: the Generation Team explores diverse query structures using large language models, the Expansion Team scales promising patterns via lightweight models, and the Management Team adaptively schedules and evaluates generation based on schema coverage and real-time query quality. This modular framework ensures a balanced trade-off between diversity, scalability, and generation cost. We apply SQL-Factory to four widely used benchmarks and generate over 300,000 SQL queries with less than \$200 API cost. Our generated queries achieve higher diversity compared to other methods, and extensive experiments demonstrate that the generated queries significantly improve the model performance in various downstream tasks.
@@ 1. Introduction
Figure 1. A conceptual illustration of SQL-Factory’s multi-agent framework. The framework consists of six agents: (1) Table Selection Agent, (2) Generation Agent, (3) Management Agent, (4) Critical Agent, (5) Seed Selection Agent and (6) Expansion Agent.
@@
Insufficient query diversity. A lack of diversity means that the generated SQL queries tend to follow similar structures, use the same types of operators, or involve limited combinations of joined tables. Template-based approaches such as TPC-DS (Nambiar and Poess, 2006) and Join Order Benchmark (JOB) (Leis et al., 2018) typically rely on fixed syntactic forms, which restrict their ability to produce varied query types. Learning-based methods such as TreeGAN (Liu et al., 2018) and LearnedSQLGen (Zhang et al., 2022) exhibit strong tendencies to overfit to frequent query skeletons in training corpora, producing workloads clustered around recurrent table subsets and operator chains. As a result, these methods fail to capture the rich variety of queries that appear in real-world workloads. Improving query diversity would help cover more edge cases and make trained models more robust in practical database environments.
@@ 2.2. SQL Generation
SQL generation aims to automatically synthesize large volumes of semantically meaningful and structurally diverse SQL queries tailored to a specific database. This task plays an essential role in supporting downstream applications such as system benchmarking, robustness testing, and data-driven model training.
@@ 2.3. LLM-based Multi-Agent System
Large language models (LLMs) have demonstrated remarkable capabilities in natural language understanding and reasoning, enabling them to perform a wide range of complex tasks with human-like proficiency. To organize these capabilities effectively, single-agent architectures typically equip an LLM with modular components such as Profile, Memory, Planning, and Action (Xi et al., 2025), which allow it to decompose tasks, retrieve relevant information, and simulate multi-step workflows. This design has proven effective in various domains, including dialogue generation, retrieval-augmented generation, and structured data interaction (Wang et al., 2024; Modarressi et al., 2023; Hu et al., 2023).
@@
Building upon this foundation, recent research has extended LLM capabilities into collaborative multi-agent systems, where multiple specialized agents coordinate to complete tasks through dynamic communication and role-based reasoning. Multi-agent systems distribute cognitive responsibilities across distinct agents, each tailored for specific subgoals or expertise areas. This paradigm fosters more interpretable, scalable, and resilient behavior, particularly in settings that demand iterative planning, task delegation, and adaptive feedback handling. Multi-agent LLM systems have shown promise in domains such as software engineering (Qian et al., 2024; Chen et al., 2023), where agents representing roles like developer, tester, and reviewer to build and refine software artifacts. Similarly, in scientific experimentation (Zheng et al., 2023) and medical diagnostics (Tang et al., 2024), these systems enable complex reasoning chains by leveraging role specialization and interactive deliberation among agents.
@@
In the context of databases, LLM-based multi-agent approaches have recently emerged as a powerful solution for tasks requiring high modularity and semantic precision. For instance, in Text-to-SQL translation, multiple agents can collaboratively handle subtasks such as schema linking, query generation, and final SQL refinement, leading to more accurate SQL outputs (Wang et al., 2025b; Xie et al., 2024). In the realm of database anomaly diagnosis, systems like D-Bot (Zhou et al., 2024) leverage large language models to automatically analyze and identify root causes of anomalies by extracting knowledge from diagnostic documents. Additionally, ROMAS (Huang et al., 2024) introduces a role-based multi-agent framework for database monitoring and planning, featuring agents assigned to specific roles (planner, monitor, and worker) collaboratively facilitate dynamic task allocation, self-monitoring, and adaptive re-planning in complex data environments. These examples collectively demonstrate the potential of multi-agent architectures to enhance reasoning modularity, improve decision accuracy, and integrate system-level feedback more effectively. Despite these advances, the application of LLM-based multi-agent systems to large-scale SQL generation remains underexplored.
@@ 3. System Overview
This section introduces SQL-Factory, a schema-aware and cost-efficient multi-agent framework for automatically synthesizing large-scale and diverse SQL query workloads. As shown in Figure 2, SQL-Factory consists of three collaborative agent teams: (1) the Management Team, (2) the Generation Team, and (3) the Expansion Team. These teams work together to address three key challenges in SQL generation, including insufficient structural diversity, lack of schema-aware planning, and high generation cost. In addition, SQL-Factory maintains a centralized SQL Pool that stores all synthesized queries throughout the process.
Generation Team. The Generation Team is designed to maximize structural diversity, tackling the problem of repetitive or overly templated query outputs. It comprises the Table Selection Agent and the Generation Agent. The Table Selection Agent performs schema-aware planning by prioritizing structurally rich or underrepresented tables using schema statistics. Then, the Generation Agent employs a powerful LLM to generate novel and semantically valid queries from scratch. These queries often include complex constructs such as nested subqueries, aggregations, and multi-table joins. By focusing on pattern exploration, the Generation Team ensures the introduction of high-diversity query structures that guide subsequent expansion.
@@
Expansion Team. The Expansion Team addresses the cost-efficiency challenge by scaling up the number of queries using lightweight generation. It includes the Seed Selection Agent and the Expansion Agent. The Seed Selection Agent identifies high-quality queries that involve under-covered tables, and the Expansion Agent uses them as seed queries to synthesize new variants via controlled augmentation. The Expansion Team concentrates on scaling promising patterns to achieve high query volume with minimal computational overhead, enabling the system to achieve both depth and breadth in SQL generation.
Management Team. The Management Team serves as the central controller of SQL-Factory, responsible for coordinating these two teams and improving overall quality by filtering out invalid or overly similar queries. It plays a crucial role in maintaining system efficiency and adaptability, directly addressing the challenge of balancing structural diversity with generation cost. This team consists of two components: the Critical Agent and the Management Agent. The Critical Agent performs real-time quality assessments using an Execution Tool to check query validity and a Retrieval Tool to estimate semantic redundancy via hybrid similarity. Based on these evaluations, the Management Agent dynamically switches between exploration (Generation Team) and exploitation (Expansion Team) to balance novelty and efficiency.
@@
In this section, we provide a detailed methodology for SQL-Factory by describing the internal design and its key components, including the Management Team, Generation Team and Expansion Teams.
@@ 4.1.1. Management Agent
(1) $S_{t+1}=\text{Agent}_{\text{M}}(S_{t},E_{t}),$
@@
Specifically, the system begins in the exploration phase, where the Generation Team synthesizes queries for underrepresented tables. After a fixed number of generation cycles, control shifts to the Expansion Team to exploit learned patterns. During this exploitation phase, the Management Agent continuously monitors evaluation information from Critical Agent. If query redundancy increases significantly or executability declines, it triggers a transition back to exploration. This alternating control mechanism allows SQL-Factory to incrementally grow a dataset that is both diverse and scalable, while preventing the generation process from collapsing into repetitive or overly common query patterns.
@@ 4.1.2. Critical Agent
Specifically, the Critical Agent comprises two modular components: the Execution Tool $\text{Tool}_{\text{E}}$, which verifies query executability, and the Retrieval Tool $\text{Tool}_{\text{R}}$, which using hybrid similarity to retrieve the most similar SQL queries. The evaluation process differs slightly depending on whether the query was produced by the Generation Team or the Expansion Team.
For SQL queries from the Generation Team, the powerful large model ensures a degree of novelty and exploratory coverage. Therefore, we focus primarily on syntactic and semantic validity. Given a set of queries $\mathcal{Q}_{t}=\{q_{t}^{1},q_{t}^{2},\dots,q_{t}^{n}\}$ at time step $t$, the Critical Agent call the Execution Tool $\text{Tool}_{\text{E}}$ to compile each query on the target database schema and reports a binary executability flag. The overall evaluation for this batch is defined as:
(2) $E_{t}=\text{Agent}_{\text{C}}(\mathcal{Q}_{t},\text{Tool}_{\text{E}}),$
@@
By integrating the above three similarity metrics, this hybrid scoring function provides a more comprehensive assessment of query similarity. It is particularly effective in identifying subtle redundancies among expanded SQL queries, and helps ensure that only genuinely novel and diverse queries are admitted into the SQL Pool.
Finally, the evaluation for the Expansion Team is defined as follows to systematically evaluate executability and diversity.
(6) $E_{t}=\text{Agent}_{\text{C}}(\mathcal{Q}_{t},\ \text{Tool}_{\text{E}},\ \text{Tool}_{\text{R}}),$
The final evaluation result $E_{t}$ is then passed to the Management Agent to support real-time scheduling decisions.
@@ 4.2.1. Table Selection Agent
(8) $\mathcal{T}^{\prime}=\text{Agent}_{\text{T}}(\mathcal{S}),\quad\text{where}\ \mathcal{T}^{\prime}\subset\mathcal{T}$
@@ 4.2.2. Generation Agent
The agent is implemented using a powerful, instruction-tuned large language model, which is prompted with both schema information for $\mathcal{T}^{\prime}$ and high-level generation instructions, as shown in Figure 4. Rather than treating each table independently, the model considers the joint schema of all selected tables and is encouraged to synthesize complex queries that include multi-table joins, nested subqueries, aggregations, and window functions. Besides, to guide the model effectively, the schema information included in the prompt is carefully constructed to ensure both syntactic validity and semantic meaningfulness of the generated queries. For each selected table, we encode not only its column names and data types, but also relevant content-aware hints to reduce the likelihood of generating vacuous or logically invalid queries. Specifically, if a column is of an enumerated type (e.g., status, category), we provide up to five representative values from the domain to help the model infer plausible filter conditions or join constraints. If the column is of a numeric or date type (e.g., int, float, date), we include its minimum and maximum values observed in the database, offering statistical context for range-based predicates (e.g., price ¿ 100). For free-text or string columns, we sample up to five typical values as textual examples, which can guide the model to generate meaningful LIKE conditions or equality comparisons. These content-aware augmentations are inserted inline with the schema description in a structured format, forming an enriched prompt that provides the model with both structural knowledge and semantic grounding. This design significantly improves the quality and contextual relevance of the generated SQL queries, particularly for tables with sparse usage history or rich attribute semantics. Finally, the output of the Generation Agent are summarized as:
@@ 4.3.1. Seed Selection Agent
To select appropriate seeds, the agent first determines a set of underutilized tables $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ by analyzing query distribution statistics via a shared statistical module $\text{Tool}_{\text{S}}$. It then search the SQL Pool to collect a small set of seed queries $\mathcal{Q}_{\text{seed}}$, each of which must reference at least one table from $\mathcal{T}^{\prime}$. These queries are selected not only based on table inclusion but also for their structural validity and potential for lexical or syntactic variation. This approach allows the Expansion Agent to operate in a pattern-aware and data-driven manner, leveraging past structures to efficiently generate new and diverse SQL variants. Formally, the the final output of Seed Selection Agent is:
(10) $\mathcal{Q}_{\text{seed}}=\text{Agent}_{\text{S}}(\mathcal{T}^{\prime})=\{q_{i}\in\mathcal{Q}_{\text{pool}}\mid\text{Tables}(q_{i})\cap\mathcal{T}^{\prime}\neq\emptyset\}.$
@@ 4.3.2. Expansion Agent
Specifically, this agent operates on the set of seed queries, which were selected based on their relevance to underrepresented tables. It uses a lightweight model to derive multiple variants based on these seed queries, employing techniques such as structural rewriting, substructure recombination. Besides, the agent is not limited to structural transformations but also introduces lexical and logical variations to ensure semantic diversity.
As shown in Figure 5, each prompt includes a small set of seed SQL queries and task-specific instructions. These instructions encourage the model to generate structurally diverse and executable variants, while inferring schema constraints implicitly from the provided queries. Formally, given a set of seed queries $\mathcal{Q}_{\text{seed}}$, the Expansion Agent generates a corresponding set of SQL variants $\mathcal{Q}_{\text{exp}}$ as follows:
@@
After expansion, the resulting queries $\mathcal{Q}_{\text{exp}}$ are passed to the Critical Agent. In this phase, queries are not only checked for executability but also evaluated for semantic redundancy using retrieval-based similarity assessment. Only queries that meet the criteria for both executability and novelty are retained and added to the SQL Pool. Through this process, the Expansion Agent supports vertical dataset expansion. It strengthens the density of high-quality query patterns within the local neighborhood of schema space, ultimately improving the dataset’s expressiveness and utility in downstream tasks. While the Generation Agent explores new structural forms, the Expansion Agent intensifies local pattern regions, ensuring both diversity and volume.